\begin{thebibliography}{25}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Deng et~al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei]{deng2009imagenet}
Jia Deng, Wei Dong, Richard Socher, Li-Jia Li, Kai Li, and Li~Fei-Fei.
\newblock Imagenet: A large-scale hierarchical image database.
\newblock In \emph{2009 IEEE conference on computer vision and pattern recognition}, pages 248--255. Ieee, 2009.

\bibitem[Dosovitskiy et~al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et~al.]{dosovitskiy2020image}
Alexey Dosovitskiy, Lucas Beyer, Alexander Kolesnikov, Dirk Weissenborn, Xiaohua Zhai, Thomas Unterthiner, Mostafa Dehghani, Matthias Minderer, Georg Heigold, Sylvain Gelly, et~al.
\newblock An image is worth 16x16 words: Transformers for image recognition at scale.
\newblock \emph{arXiv preprint arXiv:2010.11929}, 2020.

\bibitem[Dutt et~al.(2023)Dutt, Ericsson, Sanchez, Tsaftaris, and Hospedales]{dutt2023parameter}
Raman Dutt, Linus Ericsson, Pedro Sanchez, Sotirios~A Tsaftaris, and Timothy Hospedales.
\newblock Parameter-efficient fine-tuning for medical image analysis: The missed opportunity.
\newblock \emph{arXiv preprint arXiv:2305.08252}, 2023.

\bibitem[He et~al.(2021)He, Zhou, Ma, Berg-Kirkpatrick, and Neubig]{he2021towards}
Junxian He, Chunting Zhou, Xuezhe Ma, Taylor Berg-Kirkpatrick, and Graham Neubig.
\newblock Towards a unified view of parameter-efficient transfer learning.
\newblock \emph{arXiv preprint arXiv:2110.04366}, 2021.

\bibitem[He et~al.(2022)He, Chen, Xie, Li, Doll{\'a}r, and Girshick]{he2022masked}
Kaiming He, Xinlei Chen, Saining Xie, Yanghao Li, Piotr Doll{\'a}r, and Ross Girshick.
\newblock Masked autoencoders are scalable vision learners.
\newblock In \emph{Proceedings of the IEEE/CVF conference on computer vision and pattern recognition}, pages 16000--16009, 2022.

\bibitem[Houlsby et~al.(2019)Houlsby, Giurgiu, Jastrzebski, Morrone, De~Laroussilhe, Gesmundo, Attariyan, and Gelly]{houlsby2019parameter}
Neil Houlsby, Andrei Giurgiu, Stanislaw Jastrzebski, Bruna Morrone, Quentin De~Laroussilhe, Andrea Gesmundo, Mona Attariyan, and Sylvain Gelly.
\newblock Parameter-efficient transfer learning for nlp.
\newblock In \emph{International Conference on Machine Learning}, pages 2790--2799. PMLR, 2019.

\bibitem[Hu et~al.(2021)Hu, Shen, Wallis, Allen-Zhu, Li, Wang, Wang, and Chen]{hu2021lora}
Edward~J Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu~Wang, and Weizhu Chen.
\newblock Lora: Low-rank adaptation of large language models.
\newblock \emph{arXiv preprint arXiv:2106.09685}, 2021.

\bibitem[Irvin et~al.(2019)Irvin, Rajpurkar, Ko, Yu, Ciurea-Ilcus, Chute, Marklund, Haghgoo, Ball, Shpanskaya, et~al.]{irvin2019chexpert}
Jeremy Irvin, Pranav Rajpurkar, Michael Ko, Yifan Yu, Silviana Ciurea-Ilcus, Chris Chute, Henrik Marklund, Behzad Haghgoo, Robyn Ball, Katie Shpanskaya, et~al.
\newblock Chexpert: A large chest radiograph dataset with uncertainty labels and expert comparison.
\newblock In \emph{Proceedings of the AAAI conference on artificial intelligence}, volume~33, pages 590--597, 2019.

\bibitem[Jia et~al.(2022)Jia, Tang, Chen, Cardie, Belongie, Hariharan, and Lim]{jia2022visual}
Menglin Jia, Luming Tang, Bor-Chun Chen, Claire Cardie, Serge Belongie, Bharath Hariharan, and Ser-Nam Lim.
\newblock Visual prompt tuning.
\newblock In \emph{European Conference on Computer Vision}, pages 709--727. Springer, 2022.

\bibitem[Jiang et~al.(2023)Jiang, Sun, Guo, Bai, Yan, Lu, and Xu]{jiang2023anatomical}
Yankai Jiang, Mingze Sun, Heng Guo, Xiaoyu Bai, Ke~Yan, Le~Lu, and Minfeng Xu.
\newblock Anatomical invariance modeling and semantic alignment for self-supervised learning in 3d medical image analysis.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on Computer Vision}, pages 15859--15869, 2023.

\bibitem[Johnson et~al.(2019)Johnson, Pollard, Greenbaum, Lungren, Deng, Peng, Lu, Mark, Berkowitz, and Horng]{johnson2019mimic}
Alistair~EW Johnson, Tom~J Pollard, Nathaniel~R Greenbaum, Matthew~P Lungren, Chih-ying Deng, Yifan Peng, Zhiyong Lu, Roger~G Mark, Seth~J Berkowitz, and Steven Horng.
\newblock Mimic-cxr-jpg, a large publicly available database of labeled chest radiographs.
\newblock \emph{arXiv preprint arXiv:1901.07042}, 2019.

\bibitem[Liu et~al.(2022)Liu, Tam, Muqeeth, Mohta, Huang, Bansal, and Raffel]{liu2022few}
Haokun Liu, Derek Tam, Mohammed Muqeeth, Jay Mohta, Tenghao Huang, Mohit Bansal, and Colin~A Raffel.
\newblock Few-shot parameter-efficient fine-tuning is better and cheaper than in-context learning.
\newblock \emph{Advances in Neural Information Processing Systems}, 35:\penalty0 1950--1965, 2022.

\bibitem[Oquab et~al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et~al.]{oquab2023dinov2}
Maxime Oquab, Timoth{\'e}e Darcet, Th{\'e}o Moutakanni, Huy Vo, Marc Szafraniec, Vasil Khalidov, Pierre Fernandez, Daniel Haziza, Francisco Massa, Alaaeldin El-Nouby, et~al.
\newblock Dinov2: Learning robust visual features without supervision.
\newblock \emph{arXiv preprint arXiv:2304.07193}, 2023.

\bibitem[Ponti et~al.(2023)Ponti, Sordoni, Bengio, and Reddy]{ponti2023combining}
Edoardo~Maria Ponti, Alessandro Sordoni, Yoshua Bengio, and Siva Reddy.
\newblock Combining parameter-efficient modules for task-level generalisation.
\newblock In \emph{Proceedings of the 17th Conference of the European Chapter of the Association for Computational Linguistics}, pages 687--702, 2023.

\bibitem[Shih et~al.(2019)Shih, Wu, Halabi, Kohli, Prevedello, Cook, Sharma, Amorosa, Arteaga, Galperin-Aizenberg, et~al.]{shih2019augmenting}
George Shih, Carol~C Wu, Safwan~S Halabi, Marc~D Kohli, Luciano~M Prevedello, Tessa~S Cook, Arjun Sharma, Judith~K Amorosa, Veronica Arteaga, Maya Galperin-Aizenberg, et~al.
\newblock Augmenting the national institutes of health chest radiograph dataset with expert annotations of possible pneumonia.
\newblock \emph{Radiology. Artificial intelligence}, 1\penalty0 (1), 2019.

\bibitem[Sung et~al.(2022)Sung, Cho, and Bansal]{sung2022vl}
Yi-Lin Sung, Jaemin Cho, and Mohit Bansal.
\newblock Vl-adapter: Parameter-efficient transfer learning for vision-and-language tasks.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 5227--5237, 2022.

\bibitem[Wang et~al.(2023)Wang, Bao, Dong, Bjorck, Peng, Liu, Aggarwal, Mohammed, Singhal, Som, et~al.]{wang2023image}
Wenhui Wang, Hangbo Bao, Li~Dong, Johan Bjorck, Zhiliang Peng, Qiang Liu, Kriti Aggarwal, Owais~Khan Mohammed, Saksham Singhal, Subhojit Som, et~al.
\newblock Image as a foreign language: Beit pretraining for vision and vision-language tasks.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, pages 19175--19186, 2023.

\bibitem[Wang et~al.(2017)Wang, Peng, Lu, Lu, Bagheri, and Summers]{wang2017chestx}
Xiaosong Wang, Yifan Peng, Le~Lu, Zhiyong Lu, Mohammadhadi Bagheri, and Ronald~M Summers.
\newblock Chestx-ray8: Hospital-scale chest x-ray database and benchmarks on weakly-supervised classification and localization of common thorax diseases.
\newblock In \emph{Proceedings of the IEEE conference on computer vision and pattern recognition}, pages 2097--2106, 2017.

\bibitem[Yang et~al.(2023)Yang, Zhu, Xie, Zhang, Chen, and Li]{yang2023aim}
Taojiannan Yang, Yi~Zhu, Yusheng Xie, Aston Zhang, Chen Chen, and Mu~Li.
\newblock Aim: Adapting image models for efficient video action recognition.
\newblock \emph{arXiv preprint arXiv:2302.03024}, 2023.

\bibitem[Yu et~al.(2020)Yu, Zhou, Ma, Bian, Chu, Liu, and Zheng]{yu2020difficulty}
Shuang Yu, Hong-Yu Zhou, Kai Ma, Cheng Bian, Chunyan Chu, Hanruo Liu, and Yefeng Zheng.
\newblock Difficulty-aware glaucoma classification with multi-rater consensus modeling.
\newblock In \emph{Medical Image Computing and Computer Assisted Intervention--MICCAI 2020: 23rd International Conference, Lima, Peru, October 4--8, 2020, Proceedings, Part I 23}, pages 741--750. Springer, 2020.

\bibitem[Zhang et~al.(2023)Zhang, Chen, Bukharin, He, Cheng, Chen, and Zhao]{zhang2023adaptive}
Qingru Zhang, Minshuo Chen, Alexander Bukharin, Pengcheng He, Yu~Cheng, Weizhu Chen, and Tuo Zhao.
\newblock Adaptive budget allocation for parameter-efficient fine-tuning.
\newblock \emph{arXiv preprint arXiv:2303.10512}, 2023.

\bibitem[Zhou et~al.(2023{\natexlab{a}})Zhou, Lian, Wang, and Yu]{zhou2023advancing}
Hong-Yu Zhou, Chenyu Lian, Liansheng Wang, and Yizhou Yu.
\newblock Advancing radiograph representation learning with masked record modeling.
\newblock \emph{arXiv preprint arXiv:2301.13155}, 2023{\natexlab{a}}.

\bibitem[Zhou et~al.(2023{\natexlab{b}})Zhou, Lu, Chen, Yang, and Yu]{zhou2023unified}
Hong-Yu Zhou, Chixiang Lu, Chaoqi Chen, Sibei Yang, and Yizhou Yu.
\newblock A unified visual information preservation framework for self-supervised pre-training in medical image analysis.
\newblock \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, 2023{\natexlab{b}}.

\bibitem[Zhou et~al.(2023{\natexlab{c}})Zhou, Yu, Wang, Zhang, Gao, Pan, Shao, Lu, Zhang, and Li]{zhou2023transformer}
Hong-Yu Zhou, Yizhou Yu, Chengdi Wang, Shu Zhang, Yuanxu Gao, Jia Pan, Jun Shao, Guangming Lu, Kang Zhang, and Weimin Li.
\newblock A transformer-based representation-learning model with unified processing of multimodal input for clinical diagnostics.
\newblock \emph{Nature Biomedical Engineering}, pages 1--13, 2023{\natexlab{c}}.

\bibitem[Zhu et~al.(2023)Zhu, Shen, Zhao, Wang, Wang, Zhao, Shen, and Wang]{zhu2023melo}
Yitao Zhu, Zhenrong Shen, Zihao Zhao, Sheng Wang, Xin Wang, Xiangyu Zhao, Dinggang Shen, and Qian Wang.
\newblock Melo: Low-rank adaptation is better than fine-tuning for medical image diagnosis.
\newblock \emph{arXiv preprint arXiv:2311.08236}, 2023.

\end{thebibliography}

\documentclass{midl} 

\usepackage{mwe} 

\usepackage{graphicx}
\usepackage{multirow}
\usepackage{adjustbox}
\usepackage{array}
\usepackage{xcolor}
\usepackage{colortbl}
\usepackage{floatrow}
\usepackage{booktabs}
\newfloatcommand{capbtabbox}{table}[][\FBwidth]
\newcommand{\plusnum}[1]{\textcolor{green!70!black}{#1}}
\newcommand{\minusnum}[1]{\textcolor{red}{#1}}

\title[PEFT for Medical Vision Foundation Models]{Less Could Be Better: Parameter-efficient Fine-tuning Advances Medical Vision Foundation Models}

\midlauthor{\Name{Chenyu Lian\nametag{$^{1}$}} \Email{cylian@stu.xmu.edu.cn}\\
\addr $^{1}$ School of Informatics, Xiamen University \AND
\Name{Hong-Yu Zhou\nametag{$^{2}$}} \Email{whuzhouhongyu@gmail.com}\\
\addr $^{2}$ Department of Biomedical Informatics, Harvard University \AND
\Name{Yizhou Yu\nametag{$^{3}$}} \Email{yizhouy@acm.org}\\
\addr $^{3}$ Department of Computer Science, The University of Hong Kong \AND
\Name{Liansheng Wang\midljointauthortext{Corresponding author}\nametag{$^{1}$}} \Email{lswang@xmu.edu.cn}
}

\begin{document}

\maketitle

\begin{abstract}
Parameter-efficient fine-tuning (PEFT) that was initially developed for exploiting pre-trained large language models has recently emerged as an effective approach to perform transfer learning on computer vision tasks. However, the effectiveness of PEFT on medical vision foundation models is still unclear and remains to be explored.
As a proof of concept, we conducted a detailed empirical study on applying PEFT to chest radiography foundation models. Specifically, we delved into LoRA, a representative PEFT method, and compared it against full-parameter fine-tuning (FFT) on two self-supervised radiography foundation models across three well-established chest radiograph datasets.
Our results showed that LoRA outperformed FFT in 13 out of 18 transfer learning tasks by at most 2.9\% using fewer than 1\% tunable parameters.
Combining LoRA with foundation models, we set up new state-of-the-art on a range of data-efficient learning tasks, such as an AUROC score of 80.6\% using 1\% labeled data on NIH ChestX-ray14.
We hope this study can evoke more attention from the community in the use of PEFT for transfer learning on medical imaging tasks.
Code and models are available at \url{https://github.com/RL4M/MED-PEFT}.
\end{abstract}

\begin{keywords}
Transfer learning, Medical vision foundation models, Chest X-ray.
\end{keywords}

\section{Introduction}
%
%
Full-parameter fine-tuning (FFT) has long been recognized and adopted as a superior technique to do transfer learning~\cite{he2022masked, wang2023image, zhou2023advancing,zhou2023transformer,yu2020difficulty}.
%
%
However, foundation models usually have a large number of parameters, and fine-tuning the full model weights can be a sub-optimal choice when the downstream task only has limited annotations. This contrast deserves more attention in medical imaging tasks where annotation is often hard to access due to issues like privacy and safety and also the rare nature of certain diseases.
%
%
On the other hand, parameter-efficient fine-tuning (PEFT)~\cite{houlsby2019parameter, hu2021lora, liu2022few} was proposed to largely reduce the number of model parameters to be tuned and has been widely used in both language~\cite{he2021towards, zhang2023adaptive, ponti2023combining} and vision tasks~\cite{jia2022visual, sung2022vl, yang2023aim}.

More recently, some studies tried applying PEFT for medical image analysis~\cite{dutt2023parameter,zhu2023melo}.
However, one limitation of these work is that they only investigated ImageNet~\cite{deng2009imagenet} pre-trained models and ignored the more generalizable vision foundation models that were trained on large-scale medical data with self-supervised learning~\cite{zhou2023unified,jiang2023anatomical}.
In this paper, we focus on LoRA~\cite{hu2021lora}, a representative PEFT method, comparing it to FFT on two self-supervised radiography foundation models across three well-established chest radiograph datasets.
Experimental results indicate that in 13 out of 18 transfer learning tasks, LoRA exhibits superior performance over FFT, sometimes by notable margins.
For instance, on the NIH ChestX-ray dataset with merely 1\% labeled data, LoRA outperforms FFT by 2.9\% with only 0.3\% tunable parameters.

\begin{table}[t]
\renewcommand\arraystretch{1.1}
\centering
\vspace{-0.7cm}
\resizebox{0.8\textwidth}{!}{%
\begin{tabular}{cc|ccc|ccc|ccc}
\toprule
\multirow{2}{*}{\begin{tabular}[c]{@{}c@{}}Pre-trained\\ Models\end{tabular}} & \multirow{2}{*}{\begin{tabular}[c]{@{}c@{}}Transfer\\ Methods\end{tabular}} & \multicolumn{3}{c|}{NIH}                                                                                                                                              & \multicolumn{3}{c|}{CheXpert}                                                                                                                                          & \multicolumn{3}{c}{RSNA}                                                                                                                                                 \\ \cline{3-11} 
                                                                              &                                                                             & 1\%                                                   & 10\%                                                  & 100\%                                                 & 1\%                                                   & 10\%                                                   & 100\%                                                 & 1\%                                                   & 10\%                                                     & 100\%                                                 \\ \hline
\multirow{3}{*}{MAE} & FFT & 74.2                                                  & 82.2                                                  & 85.6                                                  & 87.3                                                  & 90.3                                                   & 91.8                                                  & 89.6                                                  & 90.5                                                     & 93.1                                                  \\ \cline{2-11} 
                     & LoRA & \begin{tabular}[c]{@{}c@{}}77.1\\ (\plusnum{+2.9})\end{tabular} & \begin{tabular}[c]{@{}c@{}}82.9\\ (\plusnum{+0.7})\end{tabular} & \begin{tabular}[c]{@{}c@{}}85.7\\ (\plusnum{+0.1})\end{tabular} & \begin{tabular}[c]{@{}c@{}}88.4\\ (\plusnum{+1.1})\end{tabular} & \begin{tabular}[c]{@{}c@{}}91.1\\ (\plusnum{+0.8})\end{tabular}  & \begin{tabular}[c]{@{}c@{}}91.1\\ (\minusnum{-0.7})\end{tabular} & \begin{tabular}[c]{@{}c@{}}89.9\\ (\plusnum{+0.3})\end{tabular} & \begin{tabular}[c]{@{}c@{}}91.9\\ (\plusnum{+1.4})\end{tabular}    & \begin{tabular}[c]{@{}c@{}}93.3\\ (\plusnum{+0.2})\end{tabular} \\ \hline
\multirow{3}{*}{MRM} & FFT & 80.1                                                  & 84.1                                                  & 85.9                                                  & 90.5                                                  & 91.5                                                   & 91.6                                                  & 91.3                                                  & 92.8                                                     & 93.3                                                  \\ \cline{2-11} 
                     & LoRA & \begin{tabular}[c]{@{}c@{}}80.6\\ \plusnum{(+0.5)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}84.0\\ \minusnum{(-0.1)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}85.8\\ \minusnum{(-0.1)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}90.7\\ \plusnum{(+0.2)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}92.0 \\ \plusnum{(+0.5)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}91.5\\ \minusnum{(-0.1)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}91.2\\ \minusnum{(-0.1)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}93.1\\ \plusnum{(+0.3)}\end{tabular} & \begin{tabular}[c]{@{}c@{}}93.5\\ \plusnum{(+0.2)}\end{tabular} \\ \bottomrule
\end{tabular}%
}
\caption{Comparison of the classification results of FFT and LoRA on {MAE} and {MRM}, while 1\%, 10\%, and 100\% denote the ratios of labeled data used for fine-tuning.}
\label{table1}
\end{table}


\section{Experiments and Analyses}
\subsection{Settings}
\textbf{Datasets.}
Three chest radiograph datasets were adopted to evaluate the performance of transfer learning, including NIH ChestX-ray (NIH)~\cite{wang2017chestx}, CheXpert\cite{irvin2019chexpert}, and RSNA pneumonia (RSNA)~\cite{shih2019augmenting}.
%
%
To analyze the data efficiency of different fine-tuning methods, we also presented results with different labeling ratios.
We employed the same data splits and evaluation metrics as of ~\cite{zhou2023advancing} except that we used the official test set instead of the validation set of CheXpert.

\noindent
\textbf{Chest Radiography Foundation Models.}
We adopted two self-supervised foundation models, {MRM}~\cite{zhou2023advancing} and MAE~\cite{he2022masked}. Both of them were pre-trained on the MIMIC-CXR~\cite{johnson2019mimic} dataset, based on which LoRA and FFT were applied and compared.

\subsection{Effectiveness of LoRA}
Table~\ref{table1} compares the classification results of FFT and LoRA based on {MAE} and {MRM}, measured by AUROC (\%).
Improvements can be observed in 13 out of 18 tasks, manifesting the universality of LoRA on different radiography foundation models and datasets.
Moreover, the outstanding performance of LoRA on 1\% and 10\% labeled data indicates its high data efficiency, which is particularly meaningful for medical imaging limited by the scarcity of data.
On 100\% labeled data, LoRA performs competitively with FFT but by tuning only 1.5\% parameters, showing the efficiency in computation and storage.

%

\subsection{Ablation Analyses}
\noindent
\textbf{LoRA Rank Analysis.}
We compare the performances of different ranks of LoRA on 1\%, 10\%, and 100\% labeled data of NIH based on {MRM}, showing that the ranks of LoRA should be increased accordingly as the data scale.
AUROC (\%) scores are reported in Table~\ref{table2}.
%

\noindent
\textbf{Pre-training Epochs Analysis.}
Pre-training was conducted on the MIMIC-CXR dataset using MAE~\cite{he2022masked} for 100 and 200 epochs.
As shown in Table~\ref{table3}, 1.2\% improvement on 1\% labeled data of NIH is observed when the pre-training epochs are extended from 100 to 200, while no improvement is witnessed for FFT.
We hypothesize that LoRA benefits from the small number of tuned parameters (0.3\%), mitigating the catastrophic forgetting.

\begin{table}[t]
\renewcommand\arraystretch{1}
\begin{floatrow}
\capbtabbox{%
    \vspace{-0.7em} 
    \resizebox{0.35\textwidth}{!}{%
        \begin{tabular}{cccc}
            \toprule
            LoRA Rank  & 2    & 4    & 8       \\ \hline
            1\% & 80.4 & \textbf{80.6} & 80.4 \\ \hline \hline

            LoRA Rank  & 8    & 16    & 32       \\ \hline
            10\% & 84.0 & \textbf{84.1} & 84.0 \\ \hline \hline

             LoRA Rank  & 16    & 32    & 64       \\ \hline
            100\% & 85.7 & \textbf{85.8} & 85.8 \\ \bottomrule
        \end{tabular}%
    }%
    \caption{LoRA ranks analysis.}
    \label{table2}
}

\renewcommand\arraystretch{1}
\capbtabbox{%
    \centering
    \vspace{-0.7em}
    \resizebox{0.5\textwidth}{!}{%
        \begin{tabular}{ccc}
            \toprule
            Methods               & Epochs of Pretraining & AUROC (\%)      \\ \hline
            \multirow{2}{*}{FFT} & 100                   & 74.4        \\ \cline{2-3} 
                                  & 200                   & 74.2 \minusnum{(-0.2)} \\ \hline
            \multirow{2}{*}{LoRA} & 100                   & 75.9        \\ \cline{2-3} 
                                  & 200                   & 77.1 \plusnum{(+1.2)} \\ \bottomrule
        \end{tabular}%
    }%
    \caption{Comparison of pre-training epochs.}
    \label{table3}
    \small
}

\end{floatrow}
\end{table}

\subsection{More Analyses on Other Vision Foundation Models}
\noindent
\textbf{Scaling up the Foundation Models.}
%
We conducted MAE pre-training using MIMIC-CXR images on ViT-Large~\cite{dosovitskiy2020image} for 200 epochs.
Table~\ref{table4} shows the further improvements when scaling up the transformer network.
It is noteworthy that the result of LoRA based on ViT-Base is even 0.9\% higher than the one of full-parameter fine-tuning on ViT-Large, and when adopting LoRA on ViT-Large, the AUROC of NIH 1\% can be further promoted to 77.7\%.

\noindent
\textbf{Fine-tuned on Natural Images Pre-trained Foundation Models.}
The results on Table~\ref{table5} show that when adopting the natural images pre-trained models Dinov2~\cite{oquab2023dinov2} by FFT, the performance is substantially below the baseline.
While showing that ChestX-ray pre-training is still necessary to ensure downstream performance, the introduction of LoRA significantly mitigates the performance gap caused by different modalities.

\begin{table}[htbp]
\begin{floatrow}
\capbtabbox{%
    \centering
    \vspace{-0.7em}
    \resizebox{0.45\textwidth}{!}{%
        \begin{tabular}{ccc}
            \toprule
            Method                & ViT Scale & AUROC (\%) \\ \hline
            \multirow{2}{*}{FFT} & Base      & 74.2  \\ \cline{2-3}
                                  & Large     & 76.2 \plusnum{(+2.0)} \\ \hline
            \multirow{2}{*}{LoRA} & Base      & 77.1 \plusnum{(+2.9)} \\ \cline{2-3}
                                  & Large     & 77.7 \plusnum{(+3.5)} \\ \bottomrule
        \end{tabular}%
    }%
    \caption{Comparison of model scales.}
    \label{table4}
}

\capbtabbox{%
    \centering
    \vspace{-0.7em}
    \resizebox{0.45\textwidth}{!}{%
        \begin{tabular}{ccc}
            \toprule            
            Method                & Model        & AUROC (\%)  \\ \hline
            \rowcolor{gray!25}
                FFT                  & MAE ViT-B16  & 74.2 \\ \hline
            \multirow{2}{*}{FFT} & Dinov2 ViT-B14 & 66.6 \minusnum{(-7.6)} \\ \cline{2-3} 
                                  & Dinov2 ViT-L14 & 70.9 \minusnum{(-3.3)} \\ \hline
            \multirow{2}{*}{LoRA} & Dinov2 ViT-B14 & 70.3 \minusnum{(-3.9)} \\ \cline{2-3} 
                                  & Dinov2 ViT-L14 & 72.5 \minusnum{(-1.7)} \\ \bottomrule
        \end{tabular}%
    }%
    \caption{On natural foundation models.}
    \label{table5}
    \small
}

\end{floatrow}
\end{table}



\bibliography{midl-samplebibliography}

\end{document}